George is 2 years older than Sam and Sam is 3 years older than Alex. The total of their
ages is 35. Write and solve the equation to find the ages of George, Sam, and Alex.

Found 2 solutions by stanbon, sunny1992:
Answer by stanbon(75887)

George is 2 years older than Sam and Sam is 3 years older than Alex.
The total of their ages is 35.
Write and solve the equation to find the ages of George, Sam, and Alex.
-------
Equations:
g = s + 2
g = a + 3
g + s + a = 35
------
Substitute for s and for a, ad solve for "g"::
g + g-2 + g-3 = 35
3g - 5 = 35
3g = 40
================
g = 40/3 = 13 1/3 yrs (George's age)
-----
s = 13 1/3 -2 = 11 1/3 yrs (Sam's age)
------
a = 13 1/3 - 3 = 10 1/3 yrs (Alex's age)
===============
Cheers,
Stan H.
===============

Answer by sunny1992(35)

let,
Alex's age = x
sam's age = x+3
george's age = x+3+2 = x+5
X+(X+3)+(X+5) = 35
3X+8=35
3X=27
X=9
ALEX AGE = 9
SAM AGE = 12
GEORGE AGE = 14